Astronomy Report – February 21st

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/21/2020

MDRS ROBOTIC OBSERVATORY Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: Tonight will be dedicated to the research project. In addition to the already active observations, NGC2403 is joining the queue.

Images submitted with this report: Attached is the final picture of M42 imaged by the MDRS-WF last night, obtained after some time learning and tweaking on Photoshop. Hope I got it quite right on this first attempt!

Problems Encountered:
– Some frames from last night had white lines crossing the field, as shown in the attached image of M101 on the Red filter. In this case, there were 4 exposures on the Red filter with the same line at the same place, which makes me believe it was not a plane or a satellite flying over us.
– A minor issue on Photoshop: I could not find the Camera raw filter menu as indicated in the Color imaging quick guide, p.12. Is there a way of making it active?


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

EVA Report – February 21st

Crew 222 EVA Report 21-02-2020
EVA # 5
Author: Nina SEDBON (Scientist)
Purpose of EVA: Change the Batteries if the LOAC and the MegaAres. Explore the white Moon site.
Start time: 9:15
End time: 11:16

Narrative: Everything went just fine. It was very cold on the rovers to go the exploration site. We changed the batteries rapidly because we gain efficiency day after day.
After that we drove to the white moon. There, we reached a little panorama on a hill. It was amazing to look quietly at the landscape. It was as if we were actually on another planet. The colorful traces of mud were very artistic. It appeased us to feel the outside, although we were wearing our flight suits.
The rovers lost their batteries quite rapidly, so we had to be careful about that.
Destination: Experiments Site. White Moon
Coordinates (use UTM NAD27 CONUS): 0518494-4250679 and 12S 517800-4254500
Participants: Nina SEDBON (Scientist) – Nicolas BOURDEAUD (HSO) – Bleuenn ROIRON (GreenHab officer) – Eishi KIM (Astronomer)
Road(s) and routes per MDRS Map: Cow Dung Road
Mode of travel: Rovers – Deimos and Spirit

Science Report – February 21st


Crew 222 Science Report

Author : Nina SEDBON

Title : How experiments are going

– LOAC : Everything works just fine. We just took the LOAC back for the next day (or days ?) because of the rain expected. We are sending the first results tomorrow.

– MegaAres : We sent the first measures and it is coherent. That is a good start.

– Music for Plants : GreenHab officer prepared the gelatin boxes for the experiment and planted the seeds. We divided the boxes into three samples : with stimulating music, with inhibiting music, without music. Because of the temperature, we will have to put the boxes in the Science Dome.

– AeroSpring : The roots starts sprouting out of the moss, it is very exciting ! We also put the same seeds into plastic boxes in order to make a comparison with gelatin based garden and moss garden. We found out that the GreenHab was actually to hot for the gelatin to solidify. We have to put them in the Science Dome.

– The LCVG : All the fixing are almost done, espcially the sewing ones that took more time than expected.

– Water Consumption : We received the screen. The system works, we are oing to put in place in the next fex days.

– Human Factors: The entire crew takes a lot of pleasures with the experiments. We are actually playing games together. It has a positive impact on the group.

Sol Summary – February 21st

Crew 222 Sol Summary Report 21-02-2020


Summary Title: Martian Philosophy

Author’s name: Benjamin Auzou

Mission Status: Work in progress

Sol Activity Summary:

In the morning, the crew woke up with matinal workout session, that was easier than the previous days for everyone.
The purpose of today’s EVA was to explore the Moons. The EVA team spends some time contemplating the landscapes and discussing on philosophy.
In the afternoon the program was a bit different from the other days, because of the need of an EVA to take the LOAC back to the Hab.
Our astronomer Eishi showed us his observations of the night.
Also our GreenHab plented the cress seeds for the music for plants experiments that will be conduct from now to the end of the crew 223’s mission.
In the end of the afternoon we spend a moment of positive psychology together like two sols ago.

Look Ahead Plan:

Tomorrow, we will finally set up the water monitoring experiment, and focus on activities in the Hab, because of the risk of rain or snow.

Anomalies in work: none

Weather: sunny

Crew Physical Status: all members safe and sound

EVA: experiments maintenance and contemplative exploration of the Moons

Reports to be filed: Journalist report, GreenHab Report, Astronomy Report, EVA Report, Photos Report, Photo of the Day, Operations Report, Science Report , EVA Request

Support Requested: Trashes in the Engineering airlock

Journalist Report – February 21st

Crew 222 Journalist Report 21Feb2020

Author: Marie Bochard, Crew Journalist

Sol 5

Title: Discussions

Today’s EVA group went all the way up North to the White Moon. The landscape was breath-taking, and it made us want to meditate while looking at the panorama. The sea of stones in front of us was alien-like. The group sat down for a few minutes enjoying a little rest after the climb up top, and soaked in the view from up the hill. We like to take everything slow here, because nothing good comes from a rush, especially when there is a risk to be hurt. The group headed back and reached the Hab in time for lunch.

During their EVA, we baked delicious bread and prepared a rice salad. The meal was marked by a very interesting discussion between all of us, on topics as vast as ecology, politics and relationships. Nicolas is an environmental activist, and we all learned a lot about what measures need to be taken to prevent a 7°C increase by 2100. It’s hard imagining what’s going on back on Earth while living confined on Mars, because none of the issues seem to matter since we are so far away and unimpacted. But these issues do matter, since problems on Earth will prevent any further exploration of the solar system, including Mars.

Being confined here, we have to find new subjects of conversation often and there is no other source of entertainment than ourselves. We did bring books and movies, but we all prefer talking and playing card games when we have free time. It seems like we’ve known each other for years, as we are all becoming closer and closer day after day.

Valentin, the crew engineer, and myself continued working on his backpack and LCVG this afternoon. LCVG stands for Liquid Cooling and Ventilation System. It is what astronauts use to cool themselves down in their spacesuits, because in space there is no air breeze to cool the suit’s atmosphere. His backpack is similar to the one here at the station, it blows air in the helmet to prevent fogging. These two systems together mimic what marsonauts will carry in their back when they will go on EVAs. The electronics were giving us trouble, but looking at every component individually helped us solve the issue. The system is almost ready to go! We plan to test it on EVA tomorrow. Once again, I’ll keep you posted.

Operations Report – February 21st

Crew 222 Operations Report 02/21/2020

SOL: 5

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Thursday generator:
Off at 5:30am
On at 9:00pm SOC was 82%

Friday generator:
Off at 6:30am
On at 9:00pm SOC was XX%

Diesel Reading – 75%

Station Propane Reading – 32% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 48 gallons

Water Meter: 149382.89 units

Water (static tank) – 367 gallons

Static to Loft Pump used – yes

Water in GreenHab – 99 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: yes
Hours: 144.8
Beginning charge: 100%
Ending charge: 69%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover yes
Hours: 132.4
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.3
Beginning charge: 100%
Ending charge: 96%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery. It’s planned to be fixed by Dr Rupert on Sunday.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com. All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 used and radio tested for the EVA. The spacesuit number 10 worked well during the EVA.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

GreenHab Report – February 21st

Crew 222 GreenHab Report – 21/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Ambient with door open

Average temperatures: 21°C

Hours of supplemental light: NA

Daily water usage for crops: 9,5 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 99 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative: Today, I has an EVA this morning so Marie, our journalist took care of the GreenHab and watered plants.
I put seeds in the gelatin i prepared yesterday, this time is is in petri boxes for our experience about UV and muscis.
Tonight i watered plants again and harvested some origan, sage, mint and some sun carrots.

Harvest: 2g sage, 2g mint, 5g origan, 60g carrots( 6 small carrots with leafs)

Support/supplies needed: N/A

Copyright © The Mars Society. All rights reserved. | Main Site